How do you solve (1/16)s^2=4 by factoring, by finding the square roots, or by using the quadratic formula?
thanks chany

Answer by mukhopadhyay(490)

(1/16)s^2=4
=> s^2 = 64
Using Factoring Technique:
=> s^2 - 64 = 0
=> (s-8)(s+8) = 0
=> s=8 or s=-8
....
Using Completing the Squares:
s^2 - 64 = 0
=>s^2 = 64
=>s=8 or s=-8
........
Using Quadratic Formula:
s^2 - 64 = 0
=> s = [-0 + sqrt(0-4*(-64)]/2 or s = [-0 - sqrt(0-4*(-64)]/2
=> s = sqrt(4*64)/2 or s = -sqrt(4*64)/2
=> s=8 or s=-8